OpenStudy (anonymous):

how do i solve this problem on a calculator? : sin^2(pi/3) - cos^2(pi/3)

OpenStudy (unklerhaukus):

`( ` `sin` `π/3` `)` `^2` `-` `(` `cos` `π/3` `)` `^2` `=`

OpenStudy (anonymous):

you do not need a calulator for this do you know what \(\sin(\frac{\pi}{3})\) is?